Problem 2

Suppose that GM's Smith estimated the following regression equation for Chevrolet automobiles:

Qc=100,000-100Pc+2,000N+50I+30P-1000P+3A+40,000P(I)

Where Qc=quantity demanded per year of Chevrolet automobiles

Pc=price of Chevrolet automobiles, in dollars

N=population of the United States, in millions

I=per capita disposable income, in dollars

Pf=price of Ford automobiles, in dollars

Pg=real price of gasoline, in cents per gallon

A=advertising expenditures by Chevrolet, in dollars per year

Pt=credit incentives to purchase Chevrolets, in percentage points below the rate of interest on borrowing in the absence of incentives

(a) indicate the change in the number of Chevrolets purchased per year (Qc) for each unite change in the independent or explanatory variables. (b) Find the value of Qc if the average value of Pc=9,000, N=200 million, I=$10,000, Pf=$8,000, Pg=80cents, and A=$200,000, and if Pi=1. (c)Derive the equation for the demand curve for Chevrolets. (d) plot it.

Starting with the estimated demand function given for Chevorlets in problem 2. Assume that the average value of the independent variables changes to N= 225 million , I=$12,000, Pf=$10,000, Pg=100 cents, A=$250,000, and Pi= 0 (i.e., the incentives are phased out. (a) Find the equation of the new demand curve for Chevorlets. (b) find the value of Qc, if Pc is $10,000.

  1. P3(a): Substitute the given value of independent variables into the equation in problem 2 and show the new equation as Qc= a – 100Pc, a is a number. See equation (4-5) as the example (p. 130).

For N, enter 225 not 225,000,000

  1. P3(b): Do not plot the demand curve, but find the value of Qc, if Pcis $10,000.
Answer #1

Qc = 100,000 - 100Pc + 2,000N + 50I + 30Pf - 1,000Pg + 3A + 40,000P(I)

P3(a)

Plugging in new values,

Qc = 100,000 - 100Pc + (2,000 x 225) + (50 x 12,000) + (30 x 10,000) - (1,000 x 100) + (3 x 250,000) + (40,000 x 0)

Qc = 100,000 - 100Pc + 450,000 + 600,000 + 300,000 - 100,000 + 750,000

Qc = 2,100,000 - 100Pc [Equation of demand curve]

P3(b)

When Pc = 10,000:

Qc = 2,100,000 - (100 x 10,000)

Qc = 2,100,000 - 1,000,000

Qc = 1,100,000